find the measure of an angle if its suppliment measures 12degrees less than three times its compliment

Answer by nerdybill(7384)

.
supplement is when two angles summed together is 180 deg.
complement is when two angles summed together is 90 deg.
.
Let x = "the angle"
then
90-x = the complement
180-x = the supplement
.
180-x = 3(90-x)-12
180-x = 270-3x-12
180-x = 258-3x
180+2x = 258
2x = 78
x = 39 degrees (measure of the angle)